### v1

@@ 7 Conclusion and Discussion
Then by Lemma D.1 there exists orthogonal matrix $\mathbf{W}$ such that
@@
Lemma D.1.
Consider the setting of Theorem 4.2. Then for $\hat{\mathbf{B}}^{(i)}$ we have
(13) $\lambda_{k}(\hat{\mathbf{B}}^{(i)})\asymp\lambda_{k}(\mathbf{B}^{(i)})\text{ for }k=1,\dots,r,$
@@
Let $\mathbf{E}_{\mathbf{B}}:=\hat{\mathbf{B}}-\mathbf{B}$ and $\delta=Tn^{-1}(n\rho_{n})^{1/2}$. Then according to Lemma D.3 we have $\|\mathbf{E}_{\mathbf{B}}\|\leq\delta$ with high probability. Then by perturbation theorem for singular values (see Problem III.6.13 in [26]) we have
@@
By Eq. (13) and Wedin’s $\sin\Theta$ Theorem (see e.g., Theorem 4.4 in Chapter 4 of [46]) we have
@@
(14) $\|\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{W}^{\top}\|\leq\|\sin\Theta(\hat{\mathbf{U}},\mathbf{U})\|^{2}\lesssim\frac{\delta^{2}}{\lambda_{r}^{2}(\mathbf{B})},$
@@
(15) $\|\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}}\|=\|-\mathbf{U}^{\top}\mathbf{E}_{\mathbf{B}}\hat{\mathbf{U}}\|\leq\|\mathbf{E}_{\mathbf{B}}\|\lesssim\delta$
with high probability. By Eq. (13), Eq. (14) and Eq. (15) we have
(16) $\|\bm{\Lambda}\mathbf{W}^{\top}-\mathbf{W}^{\top}\hat{\bm{\Lambda}}\|$ $=\|\bm{\Lambda}(\mathbf{W}^{\top}-\mathbf{U}^{\top}\hat{\mathbf{U}})+(\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}})+(\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{W}^{\top})\hat{\bm{\Lambda}}\|$
$\leq\|\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}}\|+\|\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{W}^{\top}\|\cdot(\|\bm{\Lambda}\|+\|\hat{\bm{\Lambda}}\|)$
@@
where we define $\mathbf{H}$ as an $r\times r$ matrix with entries $\mathbf{H}_{k,\ell}=(\sqrt{\lambda_{\ell}(\hat{\mathbf{B}})}+\sqrt{\lambda_{k}(\mathbf{B})})^{-1}$ for $k,\ell\in[r]$, and thus $\|\mathbf{H}\|_{\max}\lesssim\lambda_{r}^{-1/2}(\mathbf{B})$ with high probability. Therefore by Eq. (16) we have
@@
with high probability.
Lemma D.3.
Consider the setting of Theorem 4.2. For the error between $\hat{\mathbf{B}}^{(i)}$ and $\mathbf{B}^{(i)}$ we have
@@
(17) $|(\mathbf{E}_{\mathbf{D}^{\circ 2}})_{t_{1},t_{2}}|=|\hat{\mathbf{D}}_{t_{1},t_{2}}^{2}-\mathbf{D}_{t_{1},t_{2}}^{2}|\leq|\hat{\mathbf{D}}_{t_{1},t_{2}}-\mathbf{D}_{t_{1},t_{2}}|\cdot|\hat{\mathbf{D}}_{t_{1},t_{2}}+\mathbf{D}_{t_{1},t_{2}}|.$
By Theorem 4.1 we have
(18) $|\hat{\mathbf{D}}_{t_{1},t_{2}}-\mathbf{D}_{t_{1},t_{2}}|\lesssim n^{-1/2}$
with high probability. For $\mathbf{D}_{t_{1},t_{2}}$, because $|\mathbf{D}_{t_{1},t_{2}}|=n^{-1/2}\|\mathbf{X}_{t_{1}}\mathbf{W}-\mathbf{X}_{t_{2}}\|_{F}$ for some $\mathbf{W}\in\mathcal{O}_{d}$, by Eq. (10) we have
(19) $|\mathbf{D}_{t_{1},t_{2}}|\leq n^{-1/2}(\|\mathbf{X}_{t_{1}}\|_{F}+\|\mathbf{X}_{t_{2}}\|_{F})\lesssim n^{-1/2}(n\rho_{n})^{1/2}\lesssim\rho_{n}^{1/2}.$
For $\hat{\mathbf{D}}_{t_{1},t_{2}}$, by Eq. (18) and Eq. (19) we have
(20) $|\hat{\mathbf{D}}_{t_{1},t_{2}}|\leq|\mathbf{D}_{t_{1},t_{2}}|+|\hat{\mathbf{D}}_{t_{1},t_{2}}-\mathbf{D}_{t_{1},t_{2}}|\lesssim\rho_{n}^{1/2}+n^{-1/2}\lesssim\rho_{n}^{1/2}$
with high probability under the assumption $n\rho_{n}=\Omega(\log n)$. Combining Eq. (17), Eq. (18), Eq. (19) and Eq. (20) we have
@@
(21) $\|\mathbf{E}_{\mathbf{D}^{\circ 2}}\|_{\max}\lesssim n^{-1}(n\rho_{n})^{1/2}$
@@
with high probability. It follows that $\|\hat{\mathbf{B}}-\mathbf{B}\|\lesssim Tn^{-1}(n\rho_{n})^{1/2}$ with high probability.
Appendix E Proof of Theorem 4.3
For any dynamic networks $i,j\in[m]$, by Theorem 4.2 there exist $\mathbf{W}_{\mathbf{M}}^{(i)}$ and $\mathbf{W}_{\mathbf{M}}^{(j)}\in\mathcal{O}_{r}$ such that
(22) $\frac{1}{\sqrt{T}}\|\hat{\mathbf{M}}^{(i)}\mathbf{W}_{\mathbf{M}}^{(i)}-\mathbf{M}^{(i)}\|_{F}\lesssim\frac{T^{1/2}(n\rho_{n})^{1/2}}{n\lambda_{r}^{1/2}(\mathbf{B}^{(i)})},\quad\frac{1}{\sqrt{T}}\|\hat{\mathbf{M}}^{(j)}\mathbf{W}_{\mathbf{M}}^{(j)}-\mathbf{M}^{(j)}\|_{F}\lesssim\frac{T^{1/2}(n\rho_{n})^{1/2}}{n\lambda_{r}^{1/2}(\mathbf{B}^{(j)})}$
@@
$\mathbf{W}_{\mathbf{M}}^{(i,j)}:=\underset{\mathbf{O}\in\mathcal{O}_{r}}{\operatorname{argmin}}\|\mathbf{M}^{(i)}\mathbf{O}-\mathbf{M}^{(j)}\|_{F},\quad\hat{\mathbf{W}}_{\mathbf{M}}^{(i,j)}:=\underset{\mathbf{O}\in\mathcal{O}_{r}}{\operatorname{argmin}}\|\hat{\mathbf{M}}^{(i)}\mathbf{O}-\hat{\mathbf{M}}^{(j)}\|_{F}.$
Then for $|\hat{\mathbf{D}}^{\star}_{i,j}-\mathbf{D}^{\star}_{i,j}|$, with the identical analysis of Eq. (8) we have
(23) $|\hat{\mathbf{D}}^{\star}_{i,j}-\mathbf{D}^{\star}_{i,j}|$ $\leq\frac{1}{\sqrt{T}}\|\hat{\mathbf{M}}^{(i)}\mathbf{W}_{\mathbf{M}}^{(i)}-\mathbf{M}^{(i)}\|_{F}+\frac{1}{\sqrt{T}}\|\hat{\mathbf{M}}^{(j)}\mathbf{W}_{\mathbf{M}}^{(j)}-\mathbf{M}^{(j)}\|_{F}$
@@
Furthermore, with the identical analysis of Eq. (12) and by Eq. (22) we have
(24) $\|\mathbf{W}_{\mathbf{M}}^{(i,j)}-\mathbf{W}_{\mathbf{M}}^{(i)\top}\hat{\mathbf{W}}_{\mathbf{M}}^{(i,j)}\mathbf{W}_{\mathbf{M}}^{(j)}\|$ $\lesssim\big{[}\|\hat{\mathbf{M}}^{(i)}\mathbf{W}_{\mathbf{M}}^{(i)}-\mathbf{M}^{(i)}\|_{F}\cdot\|\hat{\mathbf{M}}^{(j)}\mathbf{W}_{\mathbf{M}}^{(j)}-\mathbf{M}^{(j)}\|_{F}+\|\hat{\mathbf{M}}^{(i)}\mathbf{W}_{\mathbf{M}}^{(i)}-\mathbf{M}^{(i)}\|_{F}\cdot\|\mathbf{M}^{(j)}\|_{F}$
@@
with high probability under the assumption $\frac{T(n\rho_{n})^{1/2}}{n\lambda_{r}}=O(1)$. Therefore by Eq. (22), Eq. (23) and Eq. (24) we have
@@
because $\{\mathbf{m}^{(i)}_{t}\}_{t\in[T]}$ are centered, i.e. $\sum_{t^{\prime}\in[T]}\mathbf{m}^{(i)}_{t^{\prime}}=\mathbf{0}$. Under the setting of Theorem 4.2, top $r$ eigenvalues have higher order compared with remaining eigenvalues so we have
(25) $\sum_{k=1}^{r}\lambda_{k}(\mathbf{B}^{(i)})=\|\mathbf{M}^{(i)}\|_{F}^{2}=\sum_{t=1}^{T}\|\mathbf{m}^{(i)}_{t}\|^{2}\asymp T\rho_{n}\tilde{\rho}.$
Therefore when $r$ is bounded and $\mathbf{B}^{(i)}$ has bounded condition number, we have $\lambda_{r}(\mathbf{B}^{(i)})\asymp T\rho_{n}\tilde{\rho}$ and Eq. (4) in Theorem 4.2 becomes
@@
with high probability. Notice by Eq. (25) we have $\frac{1}{\sqrt{T}}\|\mathbf{M}^{(i)}\|_{F}\asymp(\rho_{n}\tilde{\rho})^{1/2}$, and therefore there exists orthogonal matrix $\mathbf{W}^{(i)}_{\mathbf{M}}$ such that $\hat{\mathbf{M}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{M}}\xrightarrow{w.h.p.}\mathbf{M}^{(i)}$ under the assumptions $n\rho_{n}\tilde{\rho}^{2}=\omega(1)$ and $T=O(n^{\ell})$ for some $\ell>0$. $\square$
Appendix G Proof of Proposition 4.5
By Theorem 4.3 and the result $\lambda_{r}(\mathbf{B}^{(i)})\asymp T\rho_{n}\tilde{\rho}$ derived in Proposition 4.4, the desired result of $\hat{\mathbf{D}}^{\star}$ follows. $\square$
Appendix H Proof of Theorem 4.8 and Theorem 4.9
For any $t_{1},t_{2}\in[T]$, according to the proof of Theorem 6 in [8] we have
(26) $|(\hat{\mathbf{D}}^{(i)})^{\circ 2}-(\breve{\mathbf{D}}^{(i)})^{\circ 2}|_{\max}\lesssim n^{-1}(n\breve{\rho}_{n})^{1/2}\log n$
with high probability. Based on Eq. (26), with almost identical analysis of Theorem 4.2 and Theorem 4.3, the desired results are derived. $\square$

### v2

@@ B.4 Proof of Theorem 4
Then by Lemma B.1 there exists orthogonal matrix $\mathbf{W}$ such that
@@
Lemma B.1.
Consider the setting of Theorem 4. Then for $\hat{\mathbf{B}}^{(i)}$ we have
$\lambda_{k}(\hat{\mathbf{B}}^{(i)})\asymp\lambda_{k}(\mathbf{B}^{(i)})\text{ for }k=1,\dots,r,$ (B.6)
@@
Let $\mathbf{E}_{\mathbf{B}}:=\hat{\mathbf{B}}-\mathbf{B}$ and $\delta=Tn^{-1}(n\rho_{n})^{1/2}$. Then according to Lemma B.2 we have $\|\mathbf{E}_{\mathbf{B}}\|\leq\delta$ with high probability. Then by perturbation theorem for singular values (see Problem III.6.13 in Horn and Johnson (2012)) we have
@@
By Eq. (B.6) and Wedin’s $\sin\Theta$ Theorem (see e.g., Theorem 4.4 in Chapter 4 of Stewart (1990)) we have
@@
$\|\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{W}^{\top}\|\leq\|\sin\Theta(\hat{\mathbf{U}},\mathbf{U})\|^{2}\lesssim\frac{\delta^{2}}{\lambda_{r}^{2}(\mathbf{B})},$ (B.7)
@@
$\|\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}}\|=\|-\mathbf{U}^{\top}\mathbf{E}_{\mathbf{B}}\hat{\mathbf{U}}\|\leq\|\mathbf{E}_{\mathbf{B}}\|\lesssim\delta$ (B.8)
with high probability. By Eq. (B.6), Eq. (B.7) and Eq. (B.8) we have
$\|\bm{\Lambda}\mathbf{W}^{\top}-\mathbf{W}^{\top}\hat{\bm{\Lambda}}\|$ $=\|\bm{\Lambda}(\mathbf{W}^{\top}-\mathbf{U}^{\top}\hat{\mathbf{U}})+(\bm{\Lambda}\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{U}^{\top}\hat{\mathbf{U}}\hat{\bm{\Lambda}})+(\mathbf{U}^{\top}\hat{\mathbf{U}}-\mathbf{W}^{\top})\hat{\bm{\Lambda}}\|$ (B.9)
[…]
@@
where we define $\mathbf{H}$ as an $r\times r$ matrix with entries $\mathbf{H}_{k,\ell}=(\sqrt{\lambda_{\ell}(\hat{\mathbf{B}})}+\sqrt{\lambda_{k}(\mathbf{B})})^{-1}$ for $k,\ell\in[r]$, and thus $\|\mathbf{H}\|_{\max}\lesssim\lambda_{r}^{-1/2}(\mathbf{B})$ with high probability. Therefore by Eq. (B.9) we have
@@
with high probability. ∎
Lemma B.2.
Consider the setting of Theorem 4. For the error between $\hat{\mathbf{B}}^{(i)}$ and $\mathbf{B}^{(i)}$ we have
@@
$|(\mathbf{E}_{\mathbf{D}^{\circ 2}})_{t_{1},t_{2}}|=|\hat{\mathbf{D}}_{t_{1},t_{2}}^{2}-\mathbf{D}_{t_{1},t_{2}}^{2}|\leq|\hat{\mathbf{D}}_{t_{1},t_{2}}-\mathbf{D}_{t_{1},t_{2}}|\cdot|\hat{\mathbf{D}}_{t_{1},t_{2}}+\mathbf{D}_{t_{1},t_{2}}|.$ (B.10)
By Theorem 3 we have
$|\hat{\mathbf{D}}_{t_{1},t_{2}}-\mathbf{D}_{t_{1},t_{2}}|\lesssim n^{-1/2}$ (B.11)
with high probability. For $\mathbf{D}_{t_{1},t_{2}}$, because $|\mathbf{D}_{t_{1},t_{2}}|=n^{-1/2}\|\mathbf{X}_{t_{1}}\mathbf{W}-\mathbf{X}_{t_{2}}\|_{F}$ for some $\mathbf{W}\in\mathcal{O}_{d}$, by Eq. (B.3) we have
$|\mathbf{D}_{t_{1},t_{2}}|\leq n^{-1/2}(\|\mathbf{X}_{t_{1}}\|_{F}+\|\mathbf{X}_{t_{2}}\|_{F})\lesssim n^{-1/2}(n\rho_{n})^{1/2}\lesssim\rho_{n}^{1/2}.$ (B.12)
For $\hat{\mathbf{D}}_{t_{1},t_{2}}$, by Eq. (B.11) and Eq. (B.12) we have
$|\hat{\mathbf{D}}_{t_{1},t_{2}}|\leq|\mathbf{D}_{t_{1},t_{2}}|+|\hat{\mathbf{D}}_{t_{1},t_{2}}-\mathbf{D}_{t_{1},t_{2}}|\lesssim\rho_{n}^{1/2}+n^{-1/2}\lesssim\rho_{n}^{1/2}$ (B.13)
with high probability under the assumption $n\rho_{n}=\Omega(\log n)$. Combining Eq. (B.10), Eq. (B.11), Eq. (B.12) and Eq. (B.13) we have
@@
$\|\mathbf{E}_{\mathbf{D}^{\circ 2}}\|_{\max}\lesssim n^{-1}(n\rho_{n})^{1/2}$ (B.14)
@@
with high probability. It follows that $\|\hat{\mathbf{B}}-\mathbf{B}\|\lesssim Tn^{-1}(n\rho_{n})^{1/2}$ with high probability. ∎
B.5 Proof of Theorem 5
For any dynamic networks $i,j\in[m]$, by Theorem 4 there exist $\mathbf{W}_{\mathbf{M}}^{(i)}$ and $\mathbf{W}_{\mathbf{M}}^{(j)}\in\mathcal{O}_{r}$ such that
$\frac{1}{\sqrt{T}}\|\hat{\mathbf{M}}^{(i)}\mathbf{W}_{\mathbf{M}}^{(i)}-\mathbf{M}^{(i)}\|_{F}\lesssim\frac{T^{1/2}(n\rho_{n})^{1/2}}{n\lambda_{r}^{1/2}(\mathbf{B}^{(i)})},\quad\frac{1}{\sqrt{T}}\|\hat{\mathbf{M}}^{(j)}\mathbf{W}_{\mathbf{M}}^{(j)}-\mathbf{M}^{(j)}\|_{F}\lesssim\frac{T^{1/2}(n\rho_{n})^{1/2}}{n\lambda_{r}^{1/2}(\mathbf{B}^{(j)})}$ (B.15)
@@ B.5 Proof of Theorem 5
$\mathbf{W}_{\mathbf{M}}^{(i,j)}=\underset{\mathbf{O}\in\mathcal{O}_{r}}{\operatorname{argmin}}\|\mathbf{M}^{(i)}\mathbf{O}-\mathbf{M}^{(j)}\|_{F},\quad\hat{\mathbf{W}}_{\mathbf{M}}^{(i,j)}=\underset{\mathbf{O}\in\mathcal{O}_{r}}{\operatorname{argmin}}\|\hat{\mathbf{M}}^{(i)}\mathbf{O}-\hat{\mathbf{M}}^{(j)}\|_{F}.$
Then for $|\hat{\mathbf{D}}^{\star}_{i,j}-\mathbf{D}^{\star}_{i,j}|$, with the identical analysis of Eq. (B.1) we have
$|\hat{\mathbf{D}}^{\star}_{i,j}-\mathbf{D}^{\star}_{i,j}|$ $\leq\frac{1}{\sqrt{T}}\|\hat{\mathbf{M}}^{(i)}\mathbf{W}_{\mathbf{M}}^{(i)}-\mathbf{M}^{(i)}\|_{F}+\frac{1}{\sqrt{T}}\|\hat{\mathbf{M}}^{(j)}\mathbf{W}_{\mathbf{M}}^{(j)}-\mathbf{M}^{(j)}\|_{F}$ (B.16)
@@
Furthermore, with the identical analysis of Eq. (B.5) and by Eq. (B.15) we have
$\|\mathbf{W}_{\mathbf{M}}^{(i,j)}-\mathbf{W}_{\mathbf{M}}^{(i)\top}\hat{\mathbf{W}}_{\mathbf{M}}^{(i,j)}\mathbf{W}_{\mathbf{M}}^{(j)}\|$ $\lesssim\big[\|\hat{\mathbf{M}}^{(i)}\mathbf{W}_{\mathbf{M}}^{(i)}-\mathbf{M}^{(i)}\|_{F}\cdot\|\hat{\mathbf{M}}^{(j)}\mathbf{W}_{\mathbf{M}}^{(j)}-\mathbf{M}^{(j)}\|_{F}+\|\hat{\mathbf{M}}^{(i)}\mathbf{W}_{\mathbf{M}}^{(i)}-\mathbf{M}^{(i)}\|_{F}\cdot\|\mathbf{M}^{(j)}\|_{F}$ (B.17)
@@
with high probability under the assumption $\frac{T(n\rho_{n})^{1/2}}{n\lambda_{r}}=O(1)$. Therefore by Eq. (B.15), Eq. (B.16) and Eq. (B.17) we have
@@ B.6 Proof of Proposition 1
because $\{\mathbf{m}^{(i)}_{t}\}_{t\in[T]}$ are centered, i.e. $\sum_{t^{\prime}\in[T]}\mathbf{m}^{(i)}_{t^{\prime}}=\mathbf{0}$. Under the setting of Theorem 4, top $r$ eigenvalues have higher order compared with remaining eigenvalues so we have
$\sum_{k=1}^{r}\lambda_{k}(\mathbf{B}^{(i)})=\|\mathbf{M}^{(i)}\|_{F}^{2}=\sum_{t=1}^{T}\|\mathbf{m}^{(i)}_{t}\|^{2}\asymp T\rho_{n}\tilde{\rho}$. (B.18)
Therefore when $r$ is bounded and $\mathbf{B}^{(i)}$ has bounded condition number, we have $\lambda_{r}(\mathbf{B}^{(i)})\asymp T\rho_{n}\tilde{\rho}$ and Eq. (4.1) in Theorem 4 becomes
@@
with high probability. Notice by Eq. (B.18) we have $\frac{1}{\sqrt{T}}\|\mathbf{M}^{(i)}\|_{F}\asymp(\rho_{n}\tilde{\rho})^{1/2}$, and therefore there exists orthogonal matrix $\mathbf{W}^{(i)}_{\mathbf{M}}$ such that $\hat{\mathbf{M}}^{(i)}\mathbf{W}^{(i)}_{\mathbf{M}}\xrightarrow{w.h.p.}\mathbf{M}^{(i)}$ under the assumptions $n\rho_{n}\tilde{\rho}^{2}=\omega(1)$ and $T=O(n^{\ell})$ for some $\ell>0$. $\square$
B.7 Proof of Proposition 2
By Theorem 5 and the result $\lambda_{r}(\mathbf{B}^{(i)})\asymp T\rho_{n}\tilde{\rho}$ derived in Proposition 1, the desired result of $\hat{\mathbf{D}}^{\star}$ follows. $\square$
B.8 Proof of Theorem 6 and Theorem 7
For any $t_{1},t_{2}\in[T]$, according to the proof of Theorem 6 in Athreya et al. (2024) we have
$|(\hat{\mathbf{D}}^{(i)})^{\circ 2}-(\breve{\mathbf{D}}^{(i)})^{\circ 2}|_{\max}\lesssim n^{-1}(n\breve{\rho}_{n})^{1/2}\log n$ (B.19)
with high probability. Based on Eq. (B.19), with almost identical analysis of Theorem 4 and Theorem 5, the desired results are derived. $\square$